# TAMP: Token-Adaptive Layerwise Pruning in Multimodal Large Language Models

### v1

@@ Abstract
Jaewoo Lee1   Keyang Xuan2   Chanakya Ekbote3 Sandeep Polisetty4Yi R. (May) Fung5Paul Pu Liang3 University of North Carolina Chapel Hill1, University of Illinois Urbana-Champaign2, Massachusetts Institute of Technology3, University of Massachusetts Amherst4, Hong Kong University of Science and Technology5 jwlee8877@gmail.com keyangx3@illinois.edu cekbote@mit.edu spolisetty@umass.edu yrfung@cse.ust.hk ppliang@media.mit.edu
@@ 3.4 Adaptive Multimodal Input Activation
A key step in our approach is identifying core input tokens by measuring their contributions. In this work, we use the last row of the attention score matrix $\mathbf{A}$ as token contributions: $\mathbf{a}\!=\!\mathbf{A}[-1,\!:]\!\in\!\mathbb{R}^{N_{v}+N_{l}}$, which captures importance of multimodal tokens. This can guide the dynamic selection of input tokens based on the unique processing demand of each block. For example, in a layer emphasizing visual information, visual tokens have high contributions $\mathbf{a}$. Thus, more visual tokens are prioritized during the selection for that layer, ensuring that its input activation retains crucial visual features.
@@ 4.3 Further Analysis and Ablation
To further validate our intuition that MLLM pruning needs to adapt to modality-specific contributions within each block, we conduct ablation studies on different token selection strategies for input activations. Our approach, AMIA, selects core tokens based on token contribution score $\mathbf{a}$ and output token distances. We examine other selection strategies: (1) Random, which randomly selects 100 tokens, and (2) Attention, which selects tokens with above-average contribution scores. As shown in Table 4 (c), attention-based selection methods, Attention and AMIA, outperform random selection, supporting our core intuition. However, the best strategy depends on the target MLLM. This may be due to the complexity of multimodal feature spaces, suggesting that further refinements in selection methods could enhance pruning robustness. We analyze AMIA’s token selection with visualizations in Appendix D.
@@ 5 Conclusion
While TAMP shows promising performance on recent MLLMs, including LLaVA-NeXT, VideoLLaMA2, and LLaVA-OneVision, this work focuses on unstructured pruning for MLLMs. However, both the main body and the appendix reveal results consistent with recent structured pruning methods, indicating the potential of our core intuitions. Future research will investigate how our approach can be extended in this direction to enhance the applicability and efficiency of MLLM pruning techniques.
@@ C.2 Block-wise Average Sparsity
Appendix D Visualizing Multimodal Selection

### v3

@@ Abstract
Jaewoo Lee1   Keyang Xuan2   Chanakya Ekbote3 Sandeep Polisetty4Yi R. (May) Fung5Paul Pu Liang3 University of North Carolina Chapel Hill1, University of Illinois Urbana-Champaign2, Massachusetts Institute of Technology3, University of Massachusetts Amherst4, Hong Kong University of Science and Technology5 jwoolee@cs.unc.edu keyangx3@illinois.edu cekbote@mit.edu spolisetty@umass.edu yrfung@cse.ust.hk ppliang@mit.edu
@@ 3.4 Adaptive Multimodal Input Activation
A key step in our approach is identifying core input tokens by measuring their contributions. In this work, we use the last row of the attention score matrix $\mathbf{A}$ as token contributions: $\mathbf{a}\!=\!\mathbf{A}[:,\!-\!1]\!\in\!\mathbb{R}^{N_{v}+N_{l}}$, which captures importance of multimodal tokens. This can guide the dynamic selection of input tokens based on the unique processing demand of each block. For example, in a layer emphasizing visual information, visual tokens have high contributions $\mathbf{a}$. Thus, more visual tokens are prioritized during the selection for that layer, ensuring that its input activation retains crucial visual features.
@@ 4.3 Further Analysis and Ablation
To further validate our intuition that MLLM pruning needs to adapt to modality-specific contributions within each block, we conduct ablation studies on different token selection strategies for input activations. Our approach, AMIA, selects core tokens based on token contribution score $\mathbf{a}$ and output token distances. We examine other selection strategies: (1) Random, which randomly selects 100 tokens, and (2) Attention, which selects tokens with above-average contribution scores. As shown in Table 4 (c), attention-based selection methods, Attention and AMIA, outperform random selection, supporting our core intuition. However, the best strategy depends on the target MLLM. This may be due to the complexity of multimodal feature spaces, suggesting that further refinements in selection methods could enhance pruning robustness. We analyze AMIA’s token selection with visualizations in Appendix E.
Audio Video Audiovisual
Method Clotho TUT Vocal Mucho Video Ego NextQA MV MUSIC Rel. (%)
-AQA 2017 Sound music MME Schema -MC -Bench -QA
Full Model 85.6 71.2 92.4 58.9 48.7 49.3 73.3 58.4 79.4 100
Shortened LLaMA 37.9 12.4 18.3 24.6 32.3 8.1 21.3 30.0 25.7 34.3
ShortGPT 64.3 42.3 77.9 31.6 43.5 31.3 25.0 45.8 44.8 64.6
OWL 60.0 22.5 73.9 24.9 50.0 27.6 21.5 36.8 36.8 55.6
ECoFLaP 21.3 5.9 22.9 27.5 8.7 5.3 19.8 27.5 26.8 27.8
DAS (Ours) 71.0 33.5 53.2 28.9 49.3 32.2 25.0 42.6 45.8 61.3
Table 5: Comparison of block pruning techniques on the VideoLLaMA2 model with 25% sparsity ratio and estimate performance on various multimodal evaluation benchmarks. The best and the second best results are in bold and underlined, respectively.
Structural Pruning.
While primary focus is on unstructured pruning of MLLMs, we also explore the applicability of our approach to structured pruning, specifically block pruning that measures the importance of blocks and removes redundant blocks accordingly. For this, we conduct experiments using the VideoLLaMA2 model. To assess block importance, we adopt several recent baselines, including Shortened LLaMA (Kim et al., 2024), ShortGPT (Men et al., 2024), OWL, and ECoFLaP. In addition, we integrate our core component, DAS, which estimates layer or block importance based on multimodal token density distributions, into a structured pruning framework.
As shown in Table 5, DAS achieves competitive or superior performance across multiple multimodal benchmarks, demonstrating its effective extension to structured pruning scenarios. Notably, ShortGPT, which computes block importance using cosine similarity between input and output token embeddings, achieves the best overall performance. This result aligns with our approach and strengthens the insight that token distributions encode rich structural information critical for pruning. By leveraging output token density distributions, our method offers stronger generalizability and effectiveness for multimodal structured pruning, in contrast to prior unstructured pruning methods such as OWL and ECoFLAP, which rely on feature outliers or gradients for importance estimation.
@@ 5 Conclusion
While TAMP shows promising performance on recent MLLMs, including LLaVA-NeXT, VideoLLaMA2, and LLaVA-OneVision, this work mainly focuses on unstructured pruning for MLLMs. However, both the main body and the appendix reveal results consistent with recent structured pruning methods, indicating the potential of our core intuitions. Although we conducted the block pruning experiment, the performance of our approach was not particularly strong. Future research will investigate how our approach can be extended and improved in this direction to enhance the applicability and efficiency of MLLM pruning techniques.
@@ C.2 Block-wise Average Sparsity
Appendix D Robustness of TAMP
Method MME- MME- ChartQA AI2D MMMU Mathvista MMBench Rel. (%)
cognition perception
Full Model 376.8 1588.3 69.2 71.7 40.1 36.2 72.2 100
SparseGPT 313.0 1447.9 65.4 64.0 34.2 31.6 65.7 88.8
Wanda 286.7 1348.3 63.9 63.9 35.2 30.3 63.8 86.1
ECoFLaP 273.0 1439.9 65.7 66.0 35.3 30.8 66.2 87.9
OWL 271.9 1335.6 61.9 61.0 33.6 29.2 61.0 82.8
TAMP (Ours) 337.2 1461.3 64.8 64.8 35.9 32.2 66.5 90.9
Table 7: Comparison of pruning techniques on the LLaVA-NeXT model with 50% sparsity ratio. Reported values are the average over three different calibration sets. The best and the second best results are in bold and underlined, respectively.
Method MME- MME- ChartQA AI2D MMMU Mathvista MMBench Rel. (%)
cognition perception
Full Model - - - - - - - -
SparseGPT 23.1 16.9 0.12 0.79 0.75 1.41 0.90 0.62
Wanda 19.3 10.6 0.88 0.47 1.06 0.25 0.23 0.75
ECoFLaP 16.5 9.24 0.18 0.12 0.63 0.15 0.09 0.90
OWL 2.5 27.7 1.10 2.61 1.49 1.43 2.68 2.66
TAMP (Ours) 4.1 8.42 0.12 0.15 0.32 0.31 0.35 0.15
Table 8: Standard deviation of the performance of pruning techniques on the LLaVA-NeXT model with 50% sparsity ratio. Reported values are the average over three different calibration sets. The best and the second best results are in bold and underlined, respectively.
In this section, we evaluate the robustness of pruning methods under variations in the calibration set. Specifically, we conduct additional experiments on the LLaVA-NeXT model with 50% sparsity ratio using two additional randomly selected calibration sets. Including the original calibration set, as reported in Table 2, we compute the mean and standard deviation of performance across three trials.
As shown in Table 7, TAMP consistently outperforms all baseline methods, including SparseGPT, the second-best method, across most benchmarks, achieving the highest average relative performance. In Table 8, we observe that TAMP not only achieves the highest average score but also the smallest standard deviation across most benchmarks. Notably, the conservative estimate of TAMP’s performance ($90.9-0.15=90.75$) is still higher than the optimistic estimate of SparseGPT’s ($88.8+0.62=89.42$), suggesting a statistically significant margin that is unlikely to arise from random variations in calibration set selection.
Moreover, the consistently low standard deviations achieved by TAMP suggest that it is less sensitive to calibration data variability compared to other methods. These findings provide strong evidence of both the robustness and effectiveness of our proposed approach.
Appendix E Visualizing Multimodal Selection